Phosgene is a deadly gas and its use in warfare has been outlawed. The gas decomposes at elevated temperature:

COCl2(g) ⇄ CO(g) + Cl2(g)          Keq 8.3 × 10-4 at 352°C

What are the pressures at equilibrium if a 2.55 L metal container is charged with 2.00 bar of phosgene at 25°C and then heated to 352°C?Eignment FULL SCREEN PRINTER VERSION BACK NEXT Chapter 14, Question 51 Phosgene is a deadly gas and its use in warfare has be
Answer #1

(pCO​​​​​​)eq = (pCl2)eq = 4.03×10-2bar

(pCOCl2)eq = 1.96 bar

Change : COU, 9) = CO (9) + (1,9) quitial : 2.00 - Equilibrium: 2.00 -x Kee = (Pcoder (ander (Pcock) er 8.3x10 = (x) (x) (2: